Solve the following differential equation:



(7x+8y)dx+(8x-2)dy=0

1
Expert's answer
2022-03-09T14:58:06-0500

Since "\frac{\partial(7x+8y)}{\partial y}=8=\frac{\partial(8x-2)}{\partial x}," we conclude that the equation is exact, and hence there exists the function "u=u(x,y)" such that

"\frac{\partial u}{\partial x}=7x+8y, \ \frac{\partial u}{\partial y}=8x-2."

It follows that "u=\frac{7}2x^2+8xy+c(y)."

Then "\frac{\partial u}{\partial y}=8x+c'(y)=8x-2," and hence "c'(y)=-2." Therefore, "c(y)=-2y+C."

We conclude that the general solution of the differential equation is of the form:

"\frac{7}2x^2+8xy-2y=C."